OpenStudy (anonymous):

if csc (t) = 3, evaluate sin (pi - t)

OpenStudy (anonymous):

all i know is sin (t) = 1/3 lol

OpenStudy (anonymous):

sinpi-t = sint

OpenStudy (anonymous):

your answer is what you wrote because \[\sin(\pi-x)=\sin(x)\] what ishaan said